# SAMM: Sharded Automated Market Makers

### v1

@@ 7.1.1 Traders’ optimal action
{restatable}
lemlemmaoptimaltraderaction In $\Gamma_{n}(tf_{\textit{SAMM}})$, a trader wants to get $b^{\textit{BA}}$ token A when the system state is R. Then for the action which obtains all $b^{\textit{BA}}$ token A in one of the smallest pools with index $i^{\ast}$, where $a_{i^{\ast}}^{\text{BA}}(b^{\textit{BA}})\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})$, the trader has no less than the revenue of any other actions:
Proof Sketch.
7.1.2 Using smallest pools is a dominant strategy
Lemma 7.1.1 indicates that when multiple AMM pools have the same smallest amount of deposited tokens, acquiring all tokens in any one of them has the highest utility. Since the utility of a trader’s strategy is the linear combination of the utility of actions, it is a dominant strategy for the trader to randomly select one of the smallest pools to acquire all required tokens: {restatable}corlemmadominanttraderstrategy In $\Gamma_{n}(tf_{\textit{SAMM}})$, a dominant strategy of a BA trader is to randomly select one of the smallest pools to acquire all required tokens:
If $n_{\min}(\textbf{R})=n$, then all pools have the same amount of deposited tokens, and the trader randomly selects one of the $n$ pools. {restatable}cortheoremperfectparallelism In $\Gamma_{n}(tf_{\textit{SAMM}})$, the system state is R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$. If $\forall i,j,R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, then the perfect parallelism strategy
is a dominant strategy for the BA trader.
7.1.3 All dominant strategies use smallest pools
We have shown that only trading in one of the smallest pools is the dominant strategy for the trader. However, to later determine the best response of liquidity providers, we need to know whether there are other dominant strategies. We show that if the trader has a positive probability of taking the action of splitting a transaction or trading in a pool with not the smallest amount of deposited tokens, then she has strictly lower utility than randomly selecting one of the smallest pools to trade: {restatable}thmtheoremstrictsingle In $\Gamma_{n}(tf_{\textit{SAMM}})$, considering the following dominant strategy of the BA trader which randomly selects one of the smallest pools to acquire all required tokens:
@@ 7.1.3 All dominant strategies use smallest pools
Proof Sketch.
From the above theorem, all the best responses of the trader should only have a positive probability of taking an action that trades in exactly one of the smallest pools: {restatable}corlemmaoptimaltraderstrategyset Considering any best response strategy $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})$ of the BA trader, the strategy should only have a positive probability of taking an action that trades in one of the smallest pools:
7.2 Liquidity Provider Strategy and SPNE
Now we turn to the strategies of the liquidity providers and identify the SPNE of the game.
7.2.1 Scaffolding
We want the liquidity provider to fill up smaller pools to keep pools balanced. We call such an action the fillup action, where if the liquidity provider inputs tokens in a pool, then the pool is the smallest pool after this action. We denote the fillup action by $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$: {restatable}defideffillup_action The fillup action of a liquidity provider $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})=\left(\left(\hat{l}^{A}_{1},\hat{l}^{B}_{1}\right),\cdots,\left(\hat{l}^{A}_{n},\hat{l}^{B}_{n}\right)\right)$ is the action satisfying that if the liquidity provider inputs tokens in a pool, then the pool is one of the smallest pools after this action:
@@ 7.2.1 Scaffolding
We also define the strategy that only takes the fillup action as the fillup strategy: {restatable}defifillupstrategy The fillup strategy of a liquidity provider $\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B})$ is the strategy that only takes the fillup action:
@@
$\rho^{A}(\textbf{R})=\min_{1\leq i\leq n}R^{A}_{i}\,.$
{restatable}
lemlemmauniquefillup For any action of liquidity provider ${a_{lp}=\left(\left(l^{A}_{1},l^{B}_{1}\right),\cdots,\left(l^{A}_{n},l^{B}_{n}\right)\right)\in\mathcal{A}_{lp}(l^{A},l^{B})}$, if $\rho^{A}(\textbf{R}+a_{lp})\geq\rho^{A}(\textbf{R}+a_{lp}^{\textit{fill}})$, then $a_{lp}=a_{lp}^{\textit{fill}}$.
Proof Sketch.
We have shown that traders are incentivized to trade in smaller pools in Section 7.1, which incentivizes liquidity providers to input their tokens in smaller pools. Additionally, if the liquidity provider makes small pools larger, she would get more trading fees, which further incentivizes them to add liquidity to small pools:
{restatable}
lemlemmatradingfee For any two pools $i$ and $j$, if $R^{A}_{i}<R^{A}_{j}$, for any output amount $b^{\textit{BA}}$ of token B, the trading fee of $\textit{pool}_{i}$ is strictly smaller than the trading fee of $\textit{pool}_{j}$:
Proof Sketch.
7.2.2 Perfect parallelism under balanced pools
{restatable}
thmtheoremperfectparallelismandlp Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
@@ 7.2.2 Perfect parallelism under balanced pools
Proof Sketch.
Given that traders prefer trading in smaller pools while larger pools generate higher trading fees, the liquidity provider should increase their share in the smallest pools and enhance their sizes. This dual objective is optimally achieved by uniformly distributing tokens across all pools. ∎
{restatable}
corcorollaryalwaysstable Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
constitute an SPNE.
7.2.3 Convergence to Balanced Pools
We now show that even if the system reaches an unbalanced state, it converges to the perfect parallelism since the liquidity provider takes the fillup strategy. We can conclude that the fillup strategy is the only best response in all SPNE:
{restatable}
thmtheoremalwaysfillup In $\Gamma_{n}(\textit{tf}_{\textit{SAMM}})$, in all SPNE, the liquidity provider’s best response is the fillup strategy:
@@ 7.2.3 Convergence to Balanced Pools
Given any action $a_{lp}$ that is not the fillup action, we can construct a new action $a_{lp}^{\prime}$ that is strictly better than $a_{lp}$. By ensuring the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ is larger than that in $\textbf{R}+a_{lp}$, we increase trading fees garnered from each transaction. Moreover, this smallest pool is also the smallest before the action, maximizing the liquidity provider’s share. Consequently, the liquidity provider earns higher revenue under $a_{lp}^{\prime}$ than under $a_{lp}$. Thus, any strategy incorporating an action other than the fillup action is not optimal. Figure 2 illustrates an example construction of $a_{lp}^{\prime}$. ∎
7.2.4 Specific SPNE under deviation
Finally, we find a specific SPNE in $\Gamma_{n}(\textit{tf}_{SAMM})$ when pools are not balanced. In this SPNE, if there are multiple smallest pools, the trader uses the smallest pool in the last step. If there is more than one smallest pool in the last step, the trader selects the one with the smallest index. Denote by $i_{\min}(\textbf{R})$ the index of the pool with the smallest amount of deposited token A in R. When $i^{\ast}=i_{\min}(\textbf{R})$, we have
$\forall j,R^{A}_{i^{\ast}}\leq R^{A}_{j}$
$\forall j,R^{A}_{i^{\ast}}=R^{A}_{j}\Rightarrow i^{\ast}\leq j\,\,.$ (22)
If the pool with index $i_{\min}(\textbf{R})$ is the only pool, the BA trader would only trade in that pool. Then the liquidity provider only taking the fill-up action is the best response strategy: {restatable}thmtheoremfillup In $\Gamma_{n}(\textit{tf}_{SAMM})$, assume that the pool state in step $k$ is $\textbf{R}({k})$, $i^{\ast}=i_{\min}(\textbf{R}({k}))$ is the index of the pool with the smallest amount of deposited token A in $\textbf{R}({k})$. Then the trader strategy is to trade in the smallest pool in the last step if it is the smallest one, or randomly select one of the smallest pools, namely,
$\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}1,&\text{if}R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\text{ and }\\ &a^{\textit{BA}}=a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}})\\ \frac{1}{n_{\min}(\textbf{R})},&\text{if }R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\text{ and }\\ &a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases}$
and the liquidity provider’s fillup strategy:
are an SPNE in step $k$.
Proof Sketch.
8 Evaluation
@@
To evaluate the performance of SAMM we use the state-of-the-art Sui blockchain (§8.1). We find that the throughput of a single contract AMM is limited (§8.2), and the throughput can be improved by increasing the number of SAMM shards (§8.3). Our analysis shows that further improvement is possible by increasing the platform’s parallelism (§8.4). Simulation using real trading data shows that traders follow the desired behavior and that SAMM significantly improves the liquidity provider revenue with little sacrifice in trader cost (§8.5)
8.1 Experimental Setup
To evaluate the performance of SAMM, we implement it [^1] Link redacted for anonymous submission. in the Move language [9] in Sui [10], a state-of-the-art blockchain that supports parallel execution. Smart contracts in Sui are independent objects, and Sui executes transactions on different objects in parallel. We deploy a local Sui testnet consisting of $4$ validators, which maintain the consensus of the blockchain. We publish transactions, including transferring tokens, deploying AMM contracts, and executing AMM transactions, via an RPC interface. For all reported results, we use a machine with 2T memory and $256$ CPU cores. Several experiments deploying the Sui testnet on one machine and sending transactions on another machine produce similar results.
We run $100$ trader processes, implemented in Rust, which send transactions to smart contracts through the Sui Rust SDK for the RPC interface. Each trader sends transactions in a random interval following an exponential distribution with expectation $\frac{1}{\lambda}$. Note this experiment is only for performance evaluation, so traders follow the perfect parallelism strategy. We vary the overall frequency of transactions by setting different values of the individual $\lambda$ values. In each test, we set a target throughput. we first warm up the system by sending transactions for $500$ seconds and then measure actual frequencies and latencies for the following $100$ seconds. If the latency is stable within the 100 measurement seconds, we report the mean value. Otherwise, or if more than half of the transactions failed, we consider the experiment failed. The latency is always higher than 1 second due to Sui’s consensus protocol. We consider a latency greater than 3 seconds as a failure.
@@ 8.1 Experimental Setup
8.2 Single-Contract Bottleneck
@@
8.3 SAMM Evaluation
We implement SAMM by modifying the trading fee mechanism in the OmniSwap contract and deploying a varying number of shards (contracts). When a trader sends a transaction, she randomly selects a SAMM contract.
Figure 1 shows the average latency in different frequencies of transaction demands on different numbers of SAMM contracts. The latency in the case of one SAMM contract is indistinguishable from a single OmniSwap contract. In some instances, more than half of the transactions failed, which is marked as X in the graph. As the number of SAMM contracts increases, the system can process higher demand.
To quantify SAMM’s performance enhancements, we evaluate the throughput with a varying number of shards. We set a latency cap of $3$ seconds. As depicted in Figure 1, the latency escalates rapidly once it surpasses $2$ seconds. Therefore, choosing other latency caps beyond $2$ seconds does not significantly affect the results. The throughput with $n$ shards is thus the highest frequency that produced a latency lower than 3 seconds. Figure 3 shows the throughput increases almost linearly at the beginning and then converges to a bound.
@@ 8.3 SAMM Evaluation
Figure 3: Maximal throughput as a function of the number of SAMM shards.
8.4 Parallelization in the Underlying Platform
@@
8.5 Simulation of Trading Data
(a) Relative traders’ costs to the external market.
(b) Relative liquidity providers’ revenue to Uniswap.

### v2

@@ VI-C1 Trader Strategy
$n_{\min}(\textbf{R})=\left|\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\right|\,\,.$ (21)
We prove that randomly selecting a shard to trade without splitting is a dominant strategy for the trader, and it is strictly better than splitting or trading in a larger shard: {thm}[] In $\Gamma_{n}(tf_{\textit{SAMM}})$, considering the following dominant strategy of the BA trader which randomly selects one of the smallest pools to acquire all required tokens:
@@
Proof:
Due to the $c$-non-splitting property, trading in a single pool is strictly better than trading in multiple pools. Then the revenue of trading in one of the smallest pools is strictly better than that in any other pool due to the $c$-smaller-better property. ∎
If shards are balanced, which means all shards have the same amount of deposited tokens, we have $n_{\min}(\textbf{R})=n$. Then it is a dominant strategy for the trader to randomly select one of the $n$ pools to trade, as we intended: {cor}[] In $\Gamma_{n}(tf_{\textit{SAMM}})$, the system state is R=(($\textbf{R}=\left(\left(R^{A}_{1},R^{B}_{1},R^{S}_{1}\right),\cdots\left(R^{A}_{n},R^{B}_{n},R^{S}_{n}\right)\right)$. If $\forall i,j,R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, then the perfect parallelism strategy
is a dominant strategy for the BA trader.
VI-C2 Liquidity Provider Strategy
Our analysis of the trader strategy shows that if the shards are balanced, traders behave as intended. Next, we consider the strategies of liquidity providers in equilibrium. They should maintain the shard balance. Moreover, their incentives should rebalance the system state even in the face of attacks that break this balance.
We want a liquidity provider to fill up smaller pools to keep pools balanced. We call such an action the fillup action, where if the liquidity provider inputs tokens in a pool, then the pool is the smallest pool after this action. We denote the fillup action by $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$. Note that there is only one such action. {defi}[] The fillup action of a liquidity provider $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})=\left(\left(\hat{l}^{A}_{1},\hat{l}^{B}_{1}\right),\cdots,\left(\hat{l}^{A}_{n},\hat{l}^{B}_{n}\right)\right)$ is the action satisfying that if the liquidity provider inputs tokens in a pool, then the pool is one of the smallest pools after this action:
@@ VI-C2 Liquidity Provider Strategy
We also define the strategy that only takes the fillup action as the fillup strategy: {defi}[] The fillup strategy of a liquidity provider $\tau^{\textit{fill}}_{lp}(\textbf{R},l^{A},l^{B})$ is the strategy that only takes the fillup action:
@@
We first prove that, when all pools have identical sizes, the fillup action is to input tokens in all pools evenly, which is the best response of the liquidity provider:
{thm}
[] Denote by $\hat{a}_{lp}=\left(\left(\frac{1}{n}l^{A},\frac{1}{n}l^{B}\right),\cdots\right)$ the action of evenly depositing tokens in all pools. In $\Gamma_{n}(tf_{\textit{SAMM}})$, if for all $i$ and $j$ that the liquidity amounts are the same, $R^{A}_{i}=R^{A}_{j}$ and $R^{B}_{i}=R^{B}_{j}$, the liquidity provider strategy which only takes action $\hat{a}_{lp}$,
@@
Proof:
Traders prefer trading in smaller pools to reduce their costs. At the same time, fees are higher in larger pools. So the liquidity provider should increase her share in the smallest pools. This dual objective is optimally achieved by uniformly distributing tokens across all pools. ∎
The above theorem indicates that once the system reaches a balanced state, this state is stable. We now show that even if the system reaches an unbalanced state, maybe due to an attacker, it converges to a balanced state since the liquidity provider uses the fillup strategy.
We show that the fillup strategy is the only best response in all SPNE:
{thm}
[] In $\Gamma_{n}(\textit{tf}_{\textit{SAMM}})$, in all SPNE, the liquidity provider’s best response is the fillup strategy:
@@
Given any action $a_{lp}$ that is not the fillup action, we can construct a new action $a_{lp}^{\prime}$ that is strictly better than $a_{lp}$. By ensuring the smallest pool in $\textbf{R}+a_{lp}^{\prime}$ is larger than that in $\textbf{R}+a_{lp}$, we increase trading fees garnered from each transaction. Moreover, this smallest pool is also the smallest before the action, maximizing the liquidity provider’s share. Consequently, the liquidity provider earns higher revenue under $a_{lp}^{\prime}$ than under $a_{lp}$. Thus, any strategy incorporating an action other than the fillup action is not optimal. Figure 3 illustrates an example o  $a_{lp}^{\prime}$ construction. ∎
The above theorem does not prove the existence of an SPNE. We settle this by finding a specific SPNE in Appendix G.
In summary, our equilibrium analysis indicates that the system can achieve a stable state where the perfect parallelism strategy is the dominant strategy for traders. Moreover, the system demonstrates robustness against deviations, particularly due to attacks.
VII Evaluation
@@
To evaluate the performance of SAMM we use the state-of-the-art blockchains Sui and Solana (§VII-A). We find that the throughput of a single contract AMM is limited (§VII-B), and that it improves with the number of SAMM shards (§VII-C). Our analysis shows that further improvement is possible by increasing the platform’s parallelism (§VII-D).
VII-A Experimental Setup
We conduct experiments on the Sui [10] and Solana [56] blockchains, which support parallel execution. We first introduce the setup of the two blockchains separately and then describe the setup of the performance test.
VII-A1 Sui Setup
Smart contracts in Sui are independent objects, and Sui executes transactions on different objects in parallel. We implement SAMM [^2] Link redacted for anonymous submission. The same for Solana. in the Move language [9]. We deploy a local testnet consisting of $4$ validators maintaining the consensus of the blockchain, which follows the default configuration. The latency of transactions is always higher than 1 second due to Sui’s consensus protocol. We use Sui’s Rust RPC interface for issuing transactions.
@@ VII-A1 Sui Setup
VII-A2 Solana Setup
Each smart contract in Solana has an associated account, allowing transactions using different accounts to be executed in parallel. We implement SAMM in Rust and deploy it on the Solana testnet with one validator. The average latency of transactions is around $0.3$ seconds with low demand. We use Solana’s JavaScript RPC interface for issuing transactions.
We find that Solana’s performance is artificially limited. Specifically, Solana’s mainnet and default testnet limit the gas used by a single account within a block, and the total gas used in a block to four times the single-account gas limit [47]. In addition to experiments using the default settings (marked as Solana-Default), we conducted experiments where we compiled Solana with the gas limit for the whole block removed (marked as Solana-NoBlockLimit). This approach allows us to better evaluate execution bottlenecks. Additionally removing the gas limit for a single account results in unstable performance, with no useful results.
@@ VII-A2 Solana Setup
VII-A3 Performance Test Setup
For all reported results, we use a machine with 2T memory and $256$ CPU cores. Several experiments deploying the testnet on one machine and sending transactions on another machine produce similar results.
We run $50$ trader processes for Solanna and $100$ for Sui. Several experiments with more traders did not affect the results. Each trader sends transactions at random intervals following an exponential distribution with an expected frequency of $\lambda$. The traders send each transaction and wait for the transaction to be confirmed. They can send another transaction before the previous one is confirmed. Note this experiment is only for performance evaluation, so traders follow the perfect parallelism strategy. We vary the overall frequency of transactions by setting different values of the individual $\lambda$ values. In each test, we set a target throughput. We first warm up the system by sending transactions for $500$ seconds and then measure actual frequencies and latencies for the following $100$ seconds. If the latency is stable within the 100 measurement seconds, we report the mean value.
VII-B Single-Contract Bottleneck
To demonstrate the bottleneck of a single AMM, we first deploy a standard CPMM. We use the OmniSwap contract [38] in Sui and the token-swap contract from the Solana Program Library [32], which are generalizations of Uniswap v2. Figures 1 (${n=1}$, for Sui) and 2 (${n=1}$, for Solana-Default and Solana-NoBlockLimit) show the latencies of transaction processing in workloads with varying transaction frequencies using a single CPMM contract.
@@ VII-B Single-Contract Bottleneck
VII-C SAMM Evaluation
We implement SAMM by modifying the trading fee mechanism of CPMM contracts and deploying a varying number of shards (contracts). When a trader sends a transaction, she randomly selects a SAMM contract.
VII-C1 Sui Evaluation
Figure 1 shows the average latency for varying demand (tps) on different numbers of SAMM contracts. The latency in the case of one SAMM contract is indistinguishable from a single CPMM contract. In some instances, more than half of the transactions failed, which is marked with $\times$ in the graph. As the number of SAMM contracts increases, the system can process higher demand.
[…]
VII-C2 Solana Evaluation
[…]
We again set a latency cap of $3$ seconds and calculate the throughput. Figure 4 shows that the throughput of Solana-NoBlockLimit increases linearly up to 16 shards, while Solana-Default shows a linear increase up to 4 shards. Consequently, we observe a 16-fold throughput increase in Solana-NoBlockLimit and a 4-fold increase in Solana-Default, corresponding to the system’s capacity and the gas limit, respectively. When $n>4$, the throughput for Solana-Default remains almost constant at the level of $n=4$. Beyond $n>16$, the throughput of Solana-NoBlockLimit begins to decline, so results for $n>16$ are not shown.
Figure 4: Maximal throughput as a function of the number of SAMM shards in Sui.
Figure 5: Maximal throughput as a function of the number of SAMM shards in Solana.
VII-D Parallelization in the Underlying Platform
When there are $n$ concurrently operating AMM shards, each with a maximum throughput of $T_{\text{max}}$, the total maximum throughput, $T_{\text{total}}(n)$, is influenced by the fraction $P$ of the transaction that can be parallelized, according to Amdahl’s Law. This law states that the speedup ratio, $S(n)$, is the total throughput relative to a single AMM’s throughput, according to the expression $S(n)=\frac{1}{(1-P)+\frac{P}{n}}$. Consequently, the effective system throughput, $T_{\text{total}}(n)$, is calculated as $T_{\text{max}}\times S(n)$. For fully sequential systems like Ethereum, $P=0$, resulting in no throughput gain. For stability, Solana sets a gas limit that keeps throughput well below its serial bottleneck. Its throughput is thus linear in $n$, up to the artificial limit (Figure 5).
[…]
VIII Empirical Validation
We have seen that SAMM is incentive compatible in a game-theoretical model (§VI) and that it improves performance assuming participants’ desired behavior (§VII). To validate this in a real trading environment, we conduct simulations using real trading data (§VIII-A). We confirm (§VIII-B) that incentive compatibility is maintained for the vast majority of trades that are not too large, and throughput balance is maintained without our theoretical assumption of perfect arbitrageurs. Using performance values we see that individual traders’ costs are similar to a CPMM, but the larger throughput results in more liquidity provider revenue (§VIII-C). Nevertheless, this benefit is subject to a trade-off, as increasing trader fees reduces splits, thus improving overall throughput (§VIII-D).
VIII-A Simulation Setup
We simulate SAMM with the workload from publicly available CPMM trading data, namely the Ethereum Uniswap v2 USDC-ETH pool (hereinafter Uniswap), from block $12,000,000$ to $19,500,000$ (from 2021-03-08 to 2024-03-23, about 3 years). We conservatively choose the throughput of SAMM according to the performance evaluation of Sui since it gains less profits from throughput improvement than Solana. We simulate Uniswap v2 and SAMM using $1$ to $32$ shards. We test three different values of $c$, namely $0.003,0.005$, and $0.01$, and choose other parameters accordingly (Appendix I).
In the simulation, we do not add synthetic arbitrage transactions. There are some trades that have a larger than $c$ ratio of the pool size, especially when the pool sizes are small with large shard numbers. We still do not consider the gas fee in the simulation, which would only make the results better since it discourages trade splitting (See Section VI-A).
We run each simulation instance as follows. We uniformly at random select a point in Uniswap v2’s history (before 16500000, to ensure there are enough trades) as the starting point. For SAMM with $n$ shards, we evenly distribute the liquidity of the Uniswap contract at that point among the $n$ shards to establish the initial state of SAMM. We then simulate the real trades from that time point onward. Each trade in the real data is simulated as a trading demand with the required amount of tokens matching the output amount and tokens of the actual trade. To minimize costs, the trader selects the shard offering the lowest price and may split the trade into several smaller trades if this reduces her costs. We assume both the Uniswap and SAMM pools operate at maximal throughput, consistent with the throughput in our performance evaluation. We count trade splits, the number of transactions in different SAMM shards, liquidity provider revenue, and trader costs over 1 second, following a 1-second warm-up period. We repeat each simulation 100 times and calculate averages. Note that $1$ second in our simulation corresponds to several hours in the real world. We get historical prices of ETH to USDC from the Yahoo Finance webpage [55] and use it as the external market price. We calculate the revenue for liquidity providers and costs for traders in USDC, converting ETH amounts at the real-time price.